# FLARE: Fusing Language Models and Collaborative Architectures for Recommender Enhancement

### v1

@@
Abstract
@@
Hybrid recommender systems, combining item IDs and textual descriptions, offer potential for improved accuracy. However, previous work has largely focused on smaller datasets and model architectures. This paper introduces Flare (Fusing Language models and collaborative Architectures for Recommender Enhancement), a novel hybrid recommender that integrates a language model (mT5) with a collaborative filtering model (Bert4Rec) using a Perceiver network. This architecture allows Flare to effectively combine collaborative and content information for enhanced recommendations.
@@
We conduct a two-stage evaluation, first assessing Flare’s performance against established baselines on smaller datasets, where it demonstrates competitive accuracy. Subsequently, we evaluate Flare on a larger, more realistic dataset with a significantly larger item vocabulary, introducing new baselines for this setting. Finally, we showcase Flare’s inherent ability to support critiquing, enabling users to provide feedback and refine recommendations. We further leverage critiquing as an evaluation method to assess the model’s language understanding and its transferability to the recommendation task.
1 Introduction
@@
Sequential recommender systems incorporate the sequence of user interactions to recommend items grounded in the user’s historical preferences and temporal context (Donkers2017-xy; Hidasi2016-qn; sasrec; bert4rec). While early systems modeled sequences of items only, recent work have proposed hybrid recommenders that also model sequential context (e.g. textual descriptions, categories, attributes) for those items (Hidasi2016-zt; Huang2019-pu; Zhang2019-tr). These models suggest that the sequential item context is itself informative, with potential to improve general recommendation quality and handle cold start for new items.
With the recent trend in transformer-based and natural language models, more recent work has focused on recommender models that incorporate free-form text and attributes (recformer; s3rec; carca; u-bert; userllm; tiger). Adding text tends to improve performance over item ID baselines, e.g. Bert4Rec (bert4rec). However, to the best of our knowledge, the effectiveness of hybrid recommenders have been evaluated only on small datasets, specifically ones with a limited number of items. For instance, many models use subsets of the Amazon Product Reviews dataset that contains relatively small item inventory. Examples of common sub-sets include the Video Games (17k items), Office Products (27k) and Instruments (10k).
@@
To investigate the performance of hybrid models on larger datasets, this paper uses two recommenders. The first is a reimplementation of Bert4Rec to serve as an item-only baseline, and Flare, a simple test-bench hybrid recommender that extends Bert4Rec by incorporating a modern text encoder using a Perceiver (perceiver) network. This integration enables the synthesis of text based item context and item embeddings, potentially providing a richer representation for recommendations. Throughout this paper, ‘item-text’ denotes the concatenated key-value representation of all item context expressed in natural language.
When thoroughly tuned, Bert4Rec and Flare models outperform more complex hybrid models on certain datasets. Notably, we observe that on certain datasets, a tuned baseline Bert4Rec model, which uses only IDs, surpasses previously reported hybrid model results that use text. On datasets where Bert4Rec underperforms, our tuned model consistently outperforms the previously reported Bert4Rec baselines used to evaluate hybrid recommenders. These findings suggest that the contribution of adding text may yield less increased performance for next recommendation prediction than previously thought. That said, Flare, which incorporates text, outperforms our tuned Bert4Rec baseline, suggesting that item text improves metrics. In addition to the smaller datasets used by the current benchmarks, we also present new baseline results for Clothing, Shoes and Jewelry dataset from Amazon Product Reviews (376k) which involve larger vocabularies, where we observe similar trends.
@@
Finally, while it is natural to use standard ranking metrics to evaluate hybrid recommenders, we propose critiquing (Chen2012-dv; Gao2021-xo; Jannach2021-bp) as an additional methodology. For hybrid recommenders trained with textual context, we believe that evaluating recommendations with text critiques is a natural way to measure how well the model assimilates item and textual information. To evaluate models with critiquing, we simulate the user providing input for the desired recommendations based on recent recommended items. We present an evaluation methodology for applying critiquing using the Amazon Reviews dataset and establish baselines.
@@
  • We propose a conceptually simple hybrid recommender model using an LLM to encode text and a Perceiver network to produce contextual embeddings (Figure 1). These are then combined with ID embeddings for a richer representation.
  • We show that, despite its simplicity, Flare obtains competitive performance on small datasets, approaching state-of-the-art techniques in some cases and outperforming previously reported Bert4Rec baselines used to evaluate hybrid recommenders.
  • Additionally to the smaller datasets used in current benchmarks, we establish new baselines for hybrid models on large vocabulary datasets.
@@
  • We investigate critiquing as a promising methodology to assess and debug performance gains obtained as a result of adding text meta-data. We present a new critiquing-based metric and associated baselines.
2 Related Work
@@
Work in sequence modeling for recommendations have traditionally focused on RNNs (Hidasi2016-qn; Donkers2017-xy), and more recently, using transformers. For example, SASRec (sasrec) proposes using a transformer-based approach to model item sequences for recommendations, using causal self-attention. Bert4Rec (bert4rec) extends SASRec with BERT-style (bert) masked language-model training for item sequences. These seminal models serve as popular baselines in subsequent work.
@@
To improve general recommendation quality, follow-on efforts investigate hybrid models, which augment the sequence of item IDs with a corresponding sequence of item context (e.g. textual descriptions, categories, attributes). For example, HybridBERT4Rec (hybridbert4rec) extends BERT4Rec with a separate tower that models item context, using the set of users that consume that item as input. S3Rec (s3rec) proposes self-supervised pre-training tasks to align items and attributes by utilizing mutual information maximization. CaFe (li2022coarsetofine) explicitly models coarse-grained, or intent-based, sequences along with dense fine-grained sequences to better capture short-term vs long-term preferences. CARCA (carca) combines item sequences, associated context, and item attributes connected via cross-attention to better correlate between older and more recent items in the user profile. U-BERT (u-bert) trains user and content encoders during self-supervised pre-training with content-rich signals, followed by fine-tuning to focus on content-deficient domains. Some efforts explore replacing the explicit item ID representation and rely solely on a contextual representation. For example, Recformer (recformer) proposes an item representation comprising learned embeddings of text-based key-value pairs. TIGER (tiger) investigates a semantic ID representation for items, along with a T5-style (T5) training and generative approach for retrieval.
@@
As described in the next section, Flare shares similarities with some of the models above. It learns embeddings to represent each item, uses a pre-trained text encoder to encode item metadata, and models item sequence with a stack of transformer layers. Flare differs in that it does not require additional pre-training (aside from the “off-the-shelf” text encoder) nor fine-tuning.
Recent efforts explore incorporating collaborative filtering signals directly into a large-language model (LLM), motivated by the intuition that the language semantics encoded in the LLM promote item context interpretation. P5 (p5rec) represents all user–item signals, along with associated context, as natural language sequences that is input to an encoder/decoder model. LC-Rec (lc-rec) uses vector quantization for ID representation and LLM pre-training tasks to align items and language. CLLM4Rec (zhu2023collaborative) adds new user and item tokens to the LLM, along with a training head to model the collaborative filtering signal. Hua et al. (Hua_2023) studies methods to effectively add item IDs to the LLM token vocabulary. User-LLM (userllm) encodes user history into user embeddings, and incorporates them into an LLM via cross-attention. Harte et al. (Harte2023-gq) experiment with extracting item embeddings from a pretrained LLM to make recommendations directly, as well as to initialize Bert4Rec’s item embeddings. Unlike the above, Flare is focused on using encoders and does not explore the use of a language decoder to make recommendations.
We propose critiquing (surveyed in (Chen2012-dv; Gao2021-xo; Jannach2021-bp)) as a new task for evaluating hybrid models. Early critiquing systems enable the user to provide feedback on recommended items, typically at the feature-level, through (e.g.) suggested pre-designed tweaks (Burke1996-rv) or language-driven dialog systems (Shimazu2002-uq; Thompson2004-jj; Grasch2013-kw). Recent work (Wu2019-qx; Luo2020-uw; Luo2020-qh) explore systems that jointly learn latent representations for users, items, and key-phrases, the latter of which serves as critiques that can influence the recommendations through operations in the latent space. Others explore the interpretation of attributes (Nema2021-ad; Balog2021-cj) to support critiquing. We view hybrid models that associate item sequences with the associated textual context signals as natural extensions to the recent critiquable recommenders. As such, we believe hybrid recommenders offer a promising approach for building critiquing systems, and that it would be instructive to start evaluating such recommenders with critiquing.
3 Method
3.1 Task Definition
We formulate the recommendation task as follows. Assume a dataset consisting of user sequences $S=\{i_{1},i_{2},...,i_{n}\}$ containing items $i\in\mathcal{I}$. Each item $i$ is a tuple $(\mathrm{ID},T)$, where $\mathrm{ID}$ is the unique identifier of the item, and $T=\{t^{1},t^{2},...,t^{m}\}$ is a natural language description of the item consisting of text tokens $t$. Each description $T$ comprises the title, description category, and a short review summary given by the user. The task is to predict $i_{n+1}$ given $S$.
3.2 Flare Model Architecture
(a) Flare Model Architecture.
(b) Flare adapted to incorporate recommendation critiques. Critiques are textual descriptions, incorporated as a separate modality to the Perceiver, to direct the model’s predictions.
(c) Text-ID contrastive alignment loss, aligning textual embeddings with collaborative filtering embeddings. Gray squares are in-batch negative text-item embedding pairs, and yellow squares represent positive pairs.
@@ 3.2 Flare Model Architecture
Figure 1: Sequential recommender model architecture with text augmentation. Text tokens are encoded with a frozen (blue) LLM (mT5), and the associated embeddings are resampled with the Perceiver. Parameters in green boxes are learned using MLM.
@@
Flare augments the Bert4Rec architecture (bert4rec) with item context features (Figure 1). Sequential recommendation models like Bert4Rec use masked language modeling (MLM) with a cloze learning task, similar to BERT’s approach for language tokens. During training, a random subset of items in the input sequence is replaced with a special mask token $\langle mask\rangle$. The model’s objective is to predict these masked items based on their surrounding context. For example, an input sequence of items $[ID_{1},ID_{2},ID_{3},ID_{4},ID_{5}]$ could be transformed through random masking into $[ID_{1},\langle mask\rangle_{1},ID_{3},\langle mask\rangle_{2},ID_{5}]$. The labels for this masked sequence would then be $\langle mask\rangle_{1}=ID_{2}$, $\langle mask\rangle_{2}=ID_{4}$.
@@
Following the item embedding lookup, the inputs to the initial transformer layer can be represented as:
@@
$\mathcal{L}_{MLM}=\frac{1}{S^{m}}\sum_{ID_{m}\in S^{m}}P(ID_{m}=ID_{m}^{*}|S^{\prime})$ (1)
where $S^{m}$ is the set of masked items, $ID_{m}^{*}$ is the actual item corresponding to the masked item $ID_{m}$, $S^{\prime}$ is the input sequence with masking and $P(\cdot)$ represents the output of the softmax layer given the masked input.
@@
where $c_{i}$ is derived from a combination of the item embedding $e_{i}$ and the item context embedding $t_{i}$ representing item metadata. $t_{i}$ is generated using a pre-trained and frozen LLM encoder and a perceiver (perceiver) to encode the LLM output embeddings into a fixed length. The resulting output is concatenated with the item embedding and projected to the modified item embedding $c_{i}$ at the transformer input. This allows the model to learn masked item prediction with text as side information, which we show improves the performance on recommendation tasks. In our experiments, we use the mT5 Base model (mt5) to encode text. The output of the perceiver is then linearly projected to match the item embedding dimensions and then vector summed with the item embeddings.
@@
3.3 ID-Text Contrastive Learning
We implement an auxiliary contrastive loss to align item ID embeddings with item context embeddings as illustrated in Figure 1c. This approach is motivated by the observation that the less frequent, or “tail” items, which have limited representation in training data, can rely instead on their contextual representations. Conversely, frequently appearing items can benefit from collaborative filtering signals that can be learnt from training a sequential recommender. Aligning ID and context embeddings embeds both the collaborative filtering (CF) and contextual signals into the same latent space, allowing context to potentially fill in for gaps in CF. More precisely, we employ the InfoNCE loss function (infoNCE) as detailed below:
$\mathcal{L}_{c}=-\frac{1}{N}\sum^{N}_{i=1}\frac{exp(\frac{e_{i}\cdot t_{i}}{\tau}-m)}{exp(\frac{e_{i}\cdot t_{i}}{\tau}-m)+\sum^{N}_{j=1,j\neq i}exp(\frac{e_{i}\cdot t_{j}}{\tau})}$ (2)
where $\tau$ is the softmax temperature, and we also include a softmax additive margin term $m$ to each positive pair to increase the separation between positive and negative pairs, as proposed by (yang2019improving).
3.4 Critiquing

### v2

@@
Abstract.
Recent proposals in recommender systems represent items with their textual description, using a large language model. They show better results on standard benchmarks compared to an item ID-only model, such as Bert4Rec. In this work, we revisit the often-used Bert4Rec baseline and show that with further tuning, Bert4Rec significantly outperforms previously reported numbers, and in some datasets, is competitive with state-of-the-art models.
@@ Abstract.
With revised baselines for item ID-only models, this paper also establishes new competitive results for architectures that combine IDs and textual descriptions. We demonstrate this with Flare (Fusing Language models and collaborative Architectures for Recommender Enhancement). Flare is a novel hybrid sequence recommender that integrates a language model with a collaborative filtering model using a Perceiver network.
@@
Prior studies focus evaluation on datasets with limited-corpus size, but many commercially-applicable recommender systems common on the web must handle larger corpora. We evaluate Flare on a more realistic dataset with a significantly larger item vocabulary, introducing new baselines for this setting. This paper also showcases Flare’s inherent ability to support critiquing, enabling users to provide feedback and refine recommendations. We leverage critiquing as an evaluation method to assess the model’s language understanding and its transferability to the recommendation task.
Sequential and Hybrid Recommenders; Language Models
journalyear: 2025copyright: ccconference: Companion Proceedings of the ACM Web Conference 2025; April 28-May 2, 2025; Sydney, NSW, Australiabooktitle: Companion Proceedings of the ACM Web Conference 2025 (WWW Companion ’25), April 28-May 2, 2025, Sydney, NSW, Australiacopyright: noneccs: Computing methodologies Machine learningccs: Information systems Recommender systems
1. Introduction
@@
Modern web experiences depend on recommender systems to help users discover personalized and relevant content. A promising approach is a sequential recommender system, which incorporates the sequence of user interactions to recommend items grounded in the user’s historical preferences and temporal context (Donkers et al., 2017; Hidasi et al., 2016a; Kang and McAuley, 2018; Sun et al., 2019). While early systems modeled sequences of items only, recent works have proposed hybrid recommenders that also model sequential context (e.g. textual descriptions, categories, attributes) for those items (Hidasi et al., 2016b; Huang et al., 2019; Zhang et al., 2019). These models suggest that the sequential item context is itself informative, with potential to improve general recommendation quality and handle cold start for new items.
With the trend in transformer-based and natural language models, more recent work has focused on recommender models that incorporate free-form text and attributes (Li et al., 2023, 2024; Zhou et al., 2020; Rashed et al., 2022; Qiu et al., 2021; Ning et al., 2024; Rajput et al., 2023). Those papers show that using text tends to improve performance over item ID-only baselines, e.g. Bert4Rec (Sun et al., 2019), which is widely used as a baseline to justify new approaches. However, we discovered that when thoroughly tuned, Bert4Rec surpasses the performance of text-based recommenders, on certain datasets. On datasets where Bert4Rec underperforms, our tuned model consistently outperforms the previously reported Bert4Rec baselines often used to evaluate novel recommenders. These findings suggest that ID-only models are more competitive than assumed.
@@
With revised ID-only baselines, we investigate the performance of hybrid models that combine IDs and text. This paper presents Flare, a simple test-bench hybrid recommender that extends Bert4Rec by incorporating a modern text encoder using a Perceiver (Jaegle et al., 2021) network. This integration enables the synthesis of text-based item context and item embeddings, potentially providing a richer representation for recommendations. Flare outperforms state-of-the-art models on some datasets, suggesting that the combination of IDs and text unlocks additional gains.
To the best of our knowledge, the effectiveness of recommenders in the literature is mostly evaluated on small datasets, specifically ones with a limited number of items. For instance, many models use subsets of the Amazon Product Reviews dataset containing no more than around 40k, and often significantly fewer. In contrast, commercially-applicable recommenders must handle much larger item spaces. This gap in scale invites questions on whether contemporary approaches that combine IDs and rich metadata are still performant and feasible in a more realistic setting. We answer this with new baseline results for the Clothing, Shoes and Jewelry dataset from Amazon Product Reviews, which involves a larger vocabulary (376k). We observe similar performance trends as in smaller datasets. Flare also outperforms other contemporary models when fine-tuned and evaluated on this dataset.
@@
Finally, while it is natural to use standard ranking metrics to evaluate hybrid recommenders, this paper proposes critiquing (Chen and Pu, 2012; Gao et al., 2021; Jannach et al., 2021) as an additional methodology. For recommenders trained with textual context, we believe that evaluating recommendations with text critiques is a natural way to measure how well the model assimilates item and textual information. To evaluate models with critiquing, we simulate the user providing input for the desired recommendations based on recent recommended items. We present an evaluation methodology for applying critiquing using the Amazon Reviews dataset and establish baselines.
@@
  • Updated Bert4Rec baselines with thorough hyperparameter tuning, where they consistently outperform previously reported Bert4Rec runs, and in some datasets, outperform state-of-the-art models. They are valuable for reinterpreting conclusions from prior work that use Bert4Rec as a baseline, e.g. that the performance gains of adding text may not be as large as previously thought.
@@
  • Flare, a hybrid recommender model using a language model to encode text and a Perceiver network to produce contextual embeddings (Figure 1). These are then combined with ID embeddings for a richer representation. We show that Flare obtains competitive performance on standard datasets, re-affirming the value of text metadata.
  • New baselines for a larger vocabulary dataset: the Clothing, Shoes and Jewelry dataset from Amazon Product Reviews (376k), $\sim$10x larger than commonly-used datasets. We establish robust baselines on an updated Bert4Rec, fine-tuned Recformer (Li et al., 2023), and 3 model sizes of Flare, all of which serve as benchmarks for future work on large item vocabularies.
@@
  • Critiquing as a promising methodology to assess performance gains obtained from adding text metadata. We present a new critiquing-based metric and associated baselines.
This paper is organized as follows. The next section overviews related work. Section 3 describes Flare’s architecture, after which Section 4 presents our experiments and evaluation benchmarks. Finally, Section 5 outlines future work and concludes.
2. Related Work
@@
Work in sequence modeling for recommendations have traditionally focused on RNNs (Hidasi et al., 2016a; Donkers et al., 2017), and more recently, using transformers. For example, SASRec (Kang and McAuley, 2018) proposes using a transformer-based approach to model item sequences for recommendations, using causal self-attention. Bert4Rec (Sun et al., 2019) extends SASRec with BERT-style (Devlin et al., 2019) masked language-model training for item sequences. These seminal models serve as popular baselines in subsequent work.
@@
To improve general recommendation quality, follow-on efforts investigate hybrid models, which augment the sequence of item IDs with a corresponding sequence of item context (e.g. textual descriptions, categories, attributes). For example, HybridBERT4Rec (Channarong et al., 2022) extends Bert4Rec with a separate tower that models item context, using the set of users that consume that item as input. S3Rec (Zhou et al., 2020) proposes self-supervised pre-training tasks to align items and attributes by utilizing mutual information maximization. CaFe (Li et al., 2022) explicitly models coarse-grained, or intent-based, sequences along with dense fine-grained sequences to better capture short-term vs long-term preferences. CARCA (Rashed et al., 2022) combines item sequences, associated context, and item attributes connected via cross-attention to better correlate between older and more recent items in the user profile. U-BERT (Qiu et al., 2021) trains user and content encoders during self-supervised pre-training with content-rich signals, followed by fine-tuning to focus on content-deficient domains.
Some efforts remove dependence on an explicit item ID representation, relying solely on a contextual representation. For example, Recformer (Li et al., 2023) proposes an item representation comprising learned embeddings of text-based key-value pairs. TIGER (Rajput et al., 2023) investigates a semantic ID representation for items, along with a T5-style (Raffel et al., 2020) training and generative approach for retrieval. UniSRec (Hou et al., 2022) learns transferable item representations based on item description, along with sequence representations with contrastive tasks. UniSRec can also incorporate IDs in a fine-tuning step.
@@
As described in Section 3, Flare shares similarities with some of the models above. It learns embeddings to represent each item, uses a pre-trained text encoder to encode item metadata, and models item sequence with a stack of transformer layers. Flare differs in that it does not require additional pre-training (aside from the “off-the-shelf” text encoder) nor fine-tuning.
@@
Figure 1. Sequential recommender model architecture with text encoder augmentation. Text tokens are encoded with a frozen (blue) language model (mT5), and the associated embeddings are resampled with the Perceiver. Parameters in green boxes are learned using masked language modeling (MLM).
Recent efforts explore incorporating collaborative filtering signals directly into a large-language model (LLM), motivated by the intuition that the language semantics encoded in the LLM promote item context interpretation. P5 (Geng et al., 2022) represents all user–item signals, along with associated context, as natural language sequences that are input to an encoder/decoder model. LC-Rec (Zheng et al., 2024) uses vector quantization for ID representation and LLM pre-training tasks to align items and language. CLLM4Rec (Zhu et al., 2024) adds new user and item tokens to the LLM, along with a training head to model the collaborative filtering signal. Hua et al. (Hua et al., 2023) study methods to effectively add item IDs to the LLM token vocabulary. User-LLM (Ning et al., 2024) encodes user history into user embeddings and incorporates them into an LLM via cross-attention. Harte et al. (Harte et al., 2023) experiment with extracting item embeddings from a pretrained LLM to make recommendations directly, as well as to initialize Bert4Rec’s item embeddings. CALRec (Li et al., 2024) proposes a two-stage LLM finetuning approach using a combination of contrastive and language modeling losses. Unlike the above, Flare is focused on using encoders and does not explore the use of a language decoder to make recommendations.
We propose critiquing (surveyed in (Chen and Pu, 2012; Gao et al., 2021; Jannach et al., 2021)) as a new task for evaluating hybrid models. Early critiquing systems enable the user to provide feedback on recommended items, typically at the feature-level, through (e.g.) suggested pre-designed tweaks (Burke et al., 1996) or language-driven dialog systems (Shimazu, 2002; Thompson et al., 2004; Grasch et al., 2013). Recent work (Wu et al., 2019; Luo et al., 2020b, a) explores systems that jointly learn latent representations for users, items, and key-phrases, the latter of which serves as critiques that can influence the recommendations through operations in the latent space. Others explore the interpretation of attributes (Nema et al., 2021; Balog et al., 2021) to support critiquing. We view hybrid models that associate item sequences with the associated textual context signals as natural extensions to the recent critiquable recommenders. We believe hybrid recommenders offer a promising approach for building critiquing systems.
Figure 2. Flare adapted to incorporate recommendation critiques. Critiques are textual descriptions, incorporated as a separate modality to the Perceiver, to steer the model’s predictions.
3. Method
3.1. Task Definition
We formulate the recommendation task as follows. A dataset consists of user sequences, with each sequence $U=\{i_{1},i_{2},...,i_{n}\}$ containing items $i\in\mathcal{I}$. Each item $i$ is a tuple $(\mathrm{ID},T)$, where $\mathrm{ID}$ is the unique identifier of the item, and $T=\{t^{1},t^{2},...,t^{m}\}$ is a natural language description of the item consisting of text tokens $t$. Each description $T$ comprises textual content (e.g. item title, category, brand, price, etc). The task is to predict $i_{n+1}$ given $U$.
3.2. Flare Model Architecture
@@
Flare augments the Bert4Rec architecture (Sun et al., 2019) with item context features (Figure 1). Sequential recommendation models like Bert4Rec use masked language modeling (MLM) with a cloze learning task, similar to BERT’s approach for language tokens. During training, a random subset of items in the input sequence is replaced with a special mask token $\langle mask\rangle$. The model’s objective is to predict these masked items based on their surrounding context. For example, an input sequence of items $[ID_{1},ID_{2},ID_{3},ID_{4},ID_{5}]$ could be transformed through random masking into $[ID_{1}$, $\langle mask\rangle_{1}$, $ID_{3}$, $\langle mask\rangle_{2}$, $ID_{5}]$. The labels for this masked sequence would then be $\langle mask\rangle_{1}=ID_{2}$, $\langle mask\rangle_{2}=ID_{4}$.
@@
Following the item embedding lookup, the inputs to the initial transformer layer are represented as:
@@
(1) $\mathcal{L}_{MLM}=\frac{1}{|U^{m}|}\sum_{ID_{m}\in U^{m}}P(ID_{m}=ID_{m}^{*}|U^{\prime})$
where $U^{m}$ is the set of masked items, $ID_{m}^{*}$ is the actual item corresponding to the masked item $ID_{m}$, $U^{\prime}$ is the input sequence with masking and $P(\cdot)$ represents the output of the softmax layer given the masked input.
@@
where $c_{i}$ is derived from a combination of the item embedding $e_{i}$ and the item context embedding $t_{i}$ representing item metadata. $t_{i}$ is generated using a pre-trained and frozen LLM encoder and a perceiver (Jaegle et al., 2021) to encode the LLM output embeddings into a fixed length. The output of the perceiver is then linearly projected to match the item embedding dimensions and then vector summed with the item embeddings. This allows the model to learn masked item prediction with text as side information, which we show improves the performance on recommendation tasks. In our experiments, we use the mT5 Base model (Xue et al., 2021) to encode text.
@@
3.3. ID-Text Contrastive Learning
We implement an auxiliary contrastive loss to align item ID embeddings with item context embeddings. This approach is motivated by the observation that the less frequent, or “tail” items, which have limited representation in training data, can rely instead on their contextual representations. Conversely, frequently appearing items can benefit from collaborative filtering signals that may be learned from training a sequential recommender. Aligning ID and context embeddings places both the collaborative filtering (CF) and contextual signals into the same latent space, allowing context to potentially fill in for gaps in CF. More precisely, we employ the InfoNCE loss function (van den Oord et al., 2019) as detailed below:
(2) $\mathcal{L}_{c}=-\frac{1}{N}\sum^{N}_{i=1}\log\frac{\exp(\frac{e_{i}\cdot t_{i}}{\tau}-m)}{\exp(\frac{e_{i}\cdot t_{i}}{\tau}-m)+\sum^{N}_{j=1,j\neq i}\exp(\frac{e_{i}\cdot t_{j}}{\tau})}$
where $\tau$ is the softmax temperature, and we also include a softmax additive margin term $m$ to each positive pair to increase the separation between positive and negative pairs, as proposed by (Yang et al., 2019).
3.4. Critiquing